Two bicyclists start at the same time at a given point and travel in the same direction at average speeds of 8 miles per hour and 10 miles per hour. After how long are the bicyclists 5 miles apart?
Answer by checkley77(12844)

D=RT
D=8T
D=10T-5
8T=10T-5
8T-10T=-5
-2T=-5
T=-5/-2
T=2/5
T=2.5 HOURS THEY WILL BE 5 MILES APART.
PROOF:
8*2.5=10*2.5-5
20=25-5
20=20